OpenStudy (anonymous):

solve this equation please :D x+y+z=100 and 5x+500y+100z=10000 plz answer asap :)

OpenStudy (austinl):

We aren't supposed to give answers. We guide the asker towards a correct answer.

OpenStudy (anonymous):

ok just give me a clue :D

OpenStudy (anonymous):

factorization :)

OpenStudy (anonymous):

HOW ?! :(

OpenStudy (anonymous):

5X = 100 100X = 10,000 5x = 100 x = 100/5 x = 20 100x = 10000 x = 10000/100 x = 100

OpenStudy (anonymous):

@Booisme how did you do that ?!

OpenStudy (dan815):

sherlock is that you

OpenStudy (anonymous):

It'd actually be 3xyz+5x+500yz+100z=10000 ADD LIKE TERMS

OpenStudy (anonymous):

you have two equations and three unknowns there are an infinite number of answers http://www.wolframalpha.com/input/?i=x%2By%2Bz%3D100+and+5x%2B500y%2B100z%3D10000

OpenStudy (isaiah.feynman):

@satellite73 Excellent thinking! You're answer is the best here.

OpenStudy (anonymous):

Thanx a lot @satellite73 :D